# Practical Poisoning Attacks against Retrieval-Augmented Generation

### v1

@@
Abstract
Large language models (LLMs) have demonstrated impressive natural language processing abilities but face challenges such as hallucination and outdated knowledge. Retrieval-Augmented Generation (RAG) has emerged as a state-of-the-art approach to mitigate these issues. While RAG enhances LLM outputs, it remains vulnerable to poisoning attacks. Recent studies show that injecting poisoned text into the knowledge database can compromise RAG systems, but most existing attacks assume that the attacker can insert a sufficient number of poisoned texts per query to outnumber correct-answer texts in retrieval, an assumption that is often unrealistic. To address this limitation, we propose CorruptRAG, a practical poisoning attack against RAG systems in which the attacker injects only a single poisoned text, enhancing both feasibility and stealth. Extensive experiments across multiple datasets demonstrate that CorruptRAG achieves higher attack success rates compared to existing baselines.
1 Introduction
Large language models (LLMs) like GPT-3.5 (Brown et al., 2020), GPT-4 (Achiam et al., 2023), and GPT-4o (GPT, ) have shown impressive natural language processing capabilities. However, despite their strong performance across various tasks, LLMs still face challenges, particularly with hallucination, biases, and contextually inappropriate content. For example, lacking relevant knowledge can lead LLMs to generate inaccurate or misleading responses. Additionally, they may unintentionally reinforce training data biases or produce content misaligned with the intended context.
To tackle these challenges, Retrieval-Augmented Generation (RAG) (Karpukhin et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022; Thoppilan et al., 2022; Jiang et al., 2023; Salemi & Zamani, 2024; Chen et al., 2024a; Gao et al., 2023) has been introduced. RAG improves LLM output by retrieving relevant information from external knowledge sources in response to a user query. A typical RAG system includes three core components: a knowledge database, an LLM, and a retriever. The knowledge database contains a vast collection of trusted texts from sources like Wikipedia (Thakur et al., 2021), news (Soboroff et al., 2018), and academic papers (Voorhees et al., 2021). When a user submits a query, the retriever identifies and retrieves the top-$N$ relevant texts, which the LLM then uses as context to generate an accurate response.
@@ 1 Introduction
However, most existing attacks expand the threat landscape of RAG without fully considering their practicality. For instance, attacks like PoisonedRAG (Zou et al., 2025) are effective only when the number of poisoned texts exceeds that of the correct-answer texts within the top-$N$ retrieved texts per query. This constraint limits real-world applicability, as it requires careful manipulation to ensure poisoned texts outnumber correct-answer texts. This approach has two main drawbacks: (1) achieving this balance can be challenging, costly, and resource-intensive; (2) an increased presence of poisoned texts raises the risk of detection within the RAG system, reducing the attack’s stealth.
@@
Our Contributions: To bridge this gap, we introduce CorruptRAG, a practical poisoning attack against RAG systems. Unlike existing methods that rely on injecting multiple poisoned texts, CorruptRAG constrains the attacker to injecting only one poisoned text per query. This restriction enhances both the feasibility and stealth of the attack while still allowing the attacker to manipulate the knowledge database, ensuring that the LLM in RAG generates the attacker-desired response for a targeted query. However, solving this optimization problem presents significant challenges. First, the RAG retriever selects the top-$N$ most relevant texts for each query, but the discrete and non-linear nature of language introduces non-differentiable processing steps, making traditional gradient-based optimization ineffective. Additionally, performing gradient-based optimization would require the attacker to have full knowledge of the entire knowledge database, as well as access to the parameters of the retriever and LLM—information that is generally unavailable to the attacker. These constraints make designing an effective single-shot poisoning attack a complex and non-trivial task.
@@
To address this optimization challenge, we propose two variants of CorruptRAG: CorruptRAG-AS and CorruptRAG-AK, designed to craft effective and practical poisoned texts. CorruptRAG-AS draws inspiration from adversarial attack techniques by strategically constructing a poisoned text template that incorporates both the correct answer and the attacker’s targeted answer for each targeted query. This template is designed not only to counteract texts supporting the correct answer within the top-$N$ retrieved texts but also to increase the likelihood of generating the targeted answer. Building upon this, CorruptRAG-AK enhances generalizability by leveraging an LLM to refine the poisoned text generated by CorruptRAG-AS into adversarial knowledge. This adversarial knowledge extends the attack’s impact, enabling the LLM to generate the targeted answer not only for the specific targeted query but also for other related queries influenced by the adversarial knowledge.
@@
We evaluate CorruptRAG against four state-of-the-art poisoning attacks on RAG across three benchmark datasets. Our results demonstrate that CorruptRAG effectively manipulates RAG systems. Additionally, we assess its robustness against four advanced defense mechanisms, showing that CorruptRAG successfully bypasses these defenses while maintaining a high attack success rate. The key contributions of our work are as follows:
@@
  • We introduce CorruptRAG, a practical poisoning attack designed to compromise RAG systems.
@@
  • We evaluate CorruptRAG against existing poisoning attacks on three benchmark datasets under various practical settings. Extensive experiments show that CorruptRAG effectively compromises the RAG system and surpasses existing attacks in performance.
@@
  • We investigate multiple defense mechanisms and find that existing approaches are ineffective in mitigating the threat posed by CorruptRAG.
2 Preliminaries and Related Work
2.1 Retrieval-Augmented Generation (RAG)
@@
2.2 Attacks on LLMs and RAG
Attacks on LLMs aim to manipulate their outputs. Poisoning attacks (Shafahi et al., 2018; Fang et al., 2020; Steinhardt et al., 2017; Jia et al., 2021; Levine & Feizi, 2020) compromise training by injecting harmful data, corrupting model parameters. In contrast, prompt injection attacks (Liu et al., 2023; Perez & Ribeiro, 2022; Greshake et al., 2023; Deng et al., 2024) manipulate inference by embedding malicious content in inputs to induce attacker-desired responses. Recently, limited research has explored attacks on RAG systems (Shafran et al., 2024; Chaudhari et al., 2024; Zou et al., 2025; Xue et al., 2024; Cho et al., 2024). These attacks manipulate the output of RAG systems by injecting multiple poisoned texts into the knowledge database. The most relevant work to ours is by (Zou et al., 2025), in which the attacker uses an LLM to craft poisoned texts that can induce the RAG system to produce incorrect responses.
2.3 Defenses against Poisoning Attacks on LLMs and RAG Systems
Several defenses have been proposed to mitigate attacks on LLMs and RAG systems. Perplexity-based detections (Shafran et al., 2024; Chaudhari et al., 2024; Zou et al., 2025), initially designed to counter attacks on LLMs, have been adapted for use in RAG systems. These methods identify poisoned texts by calculating their perplexity, based on the observation that poisoned texts exhibit higher perplexity compared to benign ones. Rewriting (Zou et al., 2025; Hao et al., 2024; Shu et al., 2024) mitigates attacks by rephrasing queries before retrieval, reducing exposure to poisoned texts.
3 Threat Model
@@
Attacker’s knowledge: Note that a typical RAG system consists of three main components: a knowledge database, an LLM, and a retriever. We assume that the attacker does not have access to the texts within the knowledge database $\mathcal{D}$, nor knowledge of the LLM’s parameters or direct access to query it. For the retriever, we focus on a black-box setting, where the attacker cannot access or interact with the retriever’s internal parameters, reflecting a practical scenario in which the system’s inner workings are hidden from potential attackers. Furthermore, we assume that the attacker knows the correct answer for each targeted query. This assumption is practical, as the attacker can easily obtain the correct output from the RAG system by submitting the same targeted query before launching the attack.
Attacker’s capabilities: We assume the attacker can inject a small amount of poisoned text into the knowledge database $\mathcal{D}$ to ensure the LLM generates the attacker-selected response for each targeted query, compromising system reliability. This assumption is realistic and widely used in research (Shafran et al., 2024; Chaudhari et al., 2024; Zou et al., 2025), as many RAG systems draw from public, user-editable sources (e.g., Wikipedia, Reddit). Additionally, recent work (Carlini et al., 2024) demonstrates that Wikipedia pages can be practically manipulated for malicious purposes, further supporting this assumption.
Figure 1: The number of truly relevant texts among the top-5 retrieved for each query on Natural Questions dataset.
4 Our Attacks
4.1 Attacks as an Optimization Problem
@@
[…]
@@
To better understand the inherent constraints in a RAG system, we analyze the number of truly relevant texts (or texts implying correct answers) among the top-5 retrieved texts for each query in a standard, non-adversarial RAG setup. Using 50 queries from the Natural Questions (Kwiatkowski et al., 2019) dataset, we simulate a normal RAG system and use GPT-4o-mini to assess the relevance of the top-5 texts for each query. As shown in Figure 1, only few queries have 4 texts as ground-truth relevant, with most queries containing fewer than 3 ground-truth texts among the top-5. In contrast, the PoisonedRAG method inserts 5 poisoned texts per query into the knowledge database, ensuring that the number of poisoned texts surpasses the ground-truth texts. This shows that PoisonedRAG not only proves costly but also impractical, as it would cause the system to become dominated by poisoned rather than reliable information.
4.2 Approximating the Optimization Problem
@@
$p_{i}=p_{i}^{s}\oplus p_{i}^{h},$ (1)
where $\oplus$ represents the operation of concatenating texts. The sub-text $p_{i}^{s}$ is crafted to ensure that the poisoned text $p_{i}$ meets the Criterion I. Conversely, the sub-text $p_{i}^{h}$ is designed to ensure that the poisoned text $p_{i}$ fulfills the Criterion II. Since sub-text $p_{i}^{s}$ must ensure that the poisoned text $p_{i}$ appears among the top-$N$ texts retrieved, the attacker faces a challenge in a black-box setting, lacking access to the retriever’s parameters and top-$N$ texts. To address this, we assume that query $q_{i}$ is most similar to itself, meaning texts closely related to $q_{i}$ are likely in the top-$N$. Following (Zou et al., 2025; Chen et al., 2024b; Xue et al., 2024; Cheng et al., 2024), we set sub-text $p_{i}^{s}$ identical to the query $q_{i}$, i.e., $p_{i}^{s}=q_{i}$. Then we key challenge boils down to how to carefully design sub-text $p_{i}^{h}$ for a specific query $q_{i}$. In what follow, we introduce two variants of our proposed attacks, CorruptRAG-AS and CorruptRAG-AK, each crafted to construct the sub-text $p_{i}^{h}$ for a given query $q_{i}$.
4.3 CorruptRAG-AS
Our goal is to design $p_{i}^{h}$ so that the LLM consistently produces $A_{i}$ for query $q_{i}$, as described in Criterion II in Section 4.2. A straightforward method is to use an explicit adversarial instruction, like prompt injection (Perez & Ribeiro, 2022), to dictate the desired output, as shown below:
@@ 4.3 CorruptRAG-AS
However, these prompt injection attacks can be detected by defensive mechanisms (Wallace et al., 2024). Here, we apply the concept of an adversarial attack to craft $p_{i}^{h}$ for the query $q_{i}$. Since the retriever selects the top-$N$ texts, if the poisoned text $p_{i}$ is included, there will also be $N-1$ benign texts among the top-$N$, likely relevant to $q_{i}$ as they were retrieved. We strategically craft $p_{i}^{h}$ to make the LLM in the RAG system prioritize $p_{i}$ over the $N-1$ benign texts, increasing the chance of generating the attacker-desired response $A_{i}$ as the output. Let $C_{i}$ denote the correct answer for query $i$, where $C_{i}\neq A_{i}$. We further split $p_{i}^{h}$ into two sub-templates, $p_{i}^{h,\text{adv}}$ and $p_{i}^{h,\text{state}}$, satisfying the following equation:
$p_{i}^{h}=p_{i}^{h,\text{adv}}\oplus p_{i}^{h,\text{state}}.$ (2)
The sub-template $p_{i}^{h,\text{adv}}$ is designed to imply that the correct answer $C_{i}$ is outdated or incorrect, exploiting LLMs’ tendency to trust recent information. We define $p_{i}^{h,\text{adv}}$ in our paper as follows:
@@
$p_{i}=p_{i}^{s}\oplus p_{i}^{h}=q_{i}\oplus p_{i}^{h,\text{adv}}\oplus p_{i}^{h,\text{state}}.$ (3)
4.4 CorruptRAG-AK

### v2

@@
Abstract.
Large language models (LLMs) have demonstrated impressive natural language processing abilities but face challenges such as hallucination and outdated knowledge. Retrieval-Augmented Generation (RAG) has emerged as a state-of-the-art approach to mitigate these issues. While RAG enhances LLM outputs, it remains vulnerable to poisoning attacks. Recent studies show that injecting poisoned text into the knowledge database can compromise RAG systems, but most existing attacks assume that the attacker can insert a sufficient number of poisoned texts per query to outnumber correct-answer texts in retrieval, an assumption that is often unrealistic. To address this limitation, we propose CorruptRAG, a practical poisoning attack against RAG systems in which the attacker injects only a single poisoned text, enhancing both feasibility and stealth. Extensive experiments conducted on multiple large-scale datasets demonstrate that CorruptRAG achieves higher attack success rates than existing baselines.
1. Introduction
Large language models (LLMs) like GPT-3.5 (Brown et al., 2020), GPT-4 (Achiam et al., 2023), and GPT-4o (22) have shown impressive natural language processing capabilities. However, despite their strong performance across various tasks, LLMs still face challenges, particularly with hallucination, biases, and contextually inappropriate content. For example, lacking relevant knowledge can lead LLMs to generate inaccurate or misleading responses. Additionally, they may unintentionally reinforce training data biases or produce content misaligned with the intended context.
In order to tackle these challenges, Retrieval-Augmented Generation (RAG) (Karpukhin et al., 2020; Lewis et al., 2020; Borgeaud et al., 2022; Thoppilan et al., 2022; Jiang et al., 2023; Salemi and Zamani, 2024; Chen et al., 2024a; Gao et al., 2023; Liang et al., 2025; An et al., 2025; Zhang et al., 2025b) has been introduced. RAG improves LLM output by retrieving relevant information from external knowledge sources in response to a user query. A typical RAG system includes three core components: a knowledge database, an LLM, and a retriever. The knowledge database contains a vast collection of trusted texts from sources like Wikipedia (Thakur et al., 2021), news (Soboroff et al., 2018), and academic papers (Voorhees et al., 2021). When a user submits a query, the retriever identifies and retrieves the top-$N$ relevant texts, which the LLM then uses as context to generate an accurate response.
@@ 1. Introduction
However, most existing attacks expand the threat landscape of RAG without fully considering their practicality. For instance, attacks like PoisonedRAG (Zou et al., 2025) are effective only when the number of poisoned texts exceeds that of the correct-answer texts within the top-$N$ retrieved texts per query. This constraint limits real-world applicability, as it requires careful manipulation to ensure poisoned texts outnumber correct-answer texts. This approach has two main drawbacks: (1) achieving this balance can be challenging, costly, and resource-intensive; (2) an increased presence of poisoned texts raises the risk of detection, reducing the attack’s stealth.
@@
Our Contributions: To bridge this gap, we introduce CorruptRAG, a practical poisoning attack against RAG systems. Unlike existing methods that rely on injecting multiple poisoned texts, CorruptRAG constrains the attacker to injecting only one poisoned text per query. This restriction enhances both the feasibility and stealth of the attack while still allowing the attacker to manipulate the knowledge database, ensuring that the LLM in RAG generates the attacker-desired response for a targeted query. We frame our poisoning attacks as an optimization problem aimed at injecting a single poisoned text per query into the knowledge database. However, solving this optimization problem presents significant challenges. First, the RAG retriever selects the top-$N$ most relevant texts for each query, but the discrete and non-linear nature of language introduces non-differentiable processing steps, making traditional gradient-based optimization ineffective. Additionally, performing gradient-based optimization would require the attacker to possess complete knowledge of the entire knowledge database and access to the parameters of both the retriever and the LLM, information that is typically unavailable to the attacker. These constraints make designing an effective single-shot poisoning attack nontrivial.
@@
To address this optimization challenge, we propose two variants of CorruptRAG: CorruptRAG-AS and CorruptRAG-AK, designed to craft effective and practical poisoned texts. CorruptRAG-AS draws inspiration from adversarial attack techniques by strategically constructing a poisoned text template that incorporates both the correct answer and the targeted answer for each targeted query. This template is designed not only to counteract texts supporting the correct answer within the top-$N$ retrieved texts but also to increase the likelihood of generating the targeted answer. Building upon this, CorruptRAG-AK enhances generalizability by leveraging an LLM to refine poisoned texts generated by CorruptRAG-AS into adversarial knowledge. This adversarial knowledge extends the attack’s impact, enabling the LLM to generate the targeted answer not only for the specific targeted query but also for other related queries influenced by the adversarial knowledge.
@@
We compare CorruptRAG against four state-of-the-art baselines on three large-scale benchmark datasets. Our results demonstrate that CorruptRAG effectively manipulates RAG systems. Additionally, we assess its robustness against four advanced defense mechanisms, showing that CorruptRAG successfully bypasses these defenses while maintaining a high attack success rate. The key contributions of our work are as follows:
@@
  • We introduce CorruptRAG, a practical poisoning attack framework designed to compromise RAG systems.
@@
  • We compare CorruptRAG with four baselines on three large-scale datasets under various practical settings. Extensive experiments show that CorruptRAG effectively compromises the RAG system and surpasses existing attacks in performance.
@@
  • We investigate multiple defenses and find that existing approaches are ineffective in mitigating the threat posed by CorruptRAG.
2. Preliminaries and Related Work
2.1. Retrieval-Augmented Generation (RAG)
@@
2.2. Attacks on LLMs and RAG
Attacks on LLMs aim to manipulate their outputs. Poisoning attacks (Shafahi et al., 2018; Fang et al., 2020a; Steinhardt et al., 2017; Jia et al., 2021; Levine and Feizi, 2020; Fang et al., 2020b, 2018, 2021; Cao et al., 2021; Yin et al., 2024) compromise training by injecting harmful data, corrupting model parameters. In contrast, prompt injection attacks (Liu et al., 2023; Perez and Ribeiro, 2022; Greshake et al., 2023; Deng et al., 2024) manipulate inference by embedding malicious content in inputs to induce attacker-desired responses. Recently, limited research has explored attacks on RAG systems (Shafran et al., 2024; Chaudhari et al., 2024; Zou et al., 2025; Xue et al., 2024; Cho et al., 2024). These attacks manipulate the output of RAG systems by injecting multiple poisoned texts into the knowledge database. The most relevant work to ours is by (Zou et al., 2025), in which the attacker uses an LLM to craft poisoned texts that can induce the RAG system to produce incorrect responses.
Figure 1. The number of truly relevant texts among the top-5 retrieved for each query on Natural Questions dataset.
2.3. Defenses against Poisoning Attacks on LLMs and RAG Systems
A growing body of research has explored defenses to strengthen large LLMs and RAG systems against adversarial manipulation. The paraphrasing-based defense (Zou et al., 2025) mitigates poisoning attacks in RAG by rewording user queries before retrieval, effectively disrupting the association between attacker-crafted triggers and the targeted queries. The instructional prevention defense (Liu et al., 2024) addresses prompt injection attacks in LLM-integrated applications (Jain et al., 2023; Alon and Kamfonas, 2023; Gonen et al., 2022) by redesigning system prompts to explicitly direct the model to ignore potentially malicious or conflicting instructions embedded in user inputs. The LLM-based detection defense (Liu et al., 2024; Armstrong and Gorman, 2022; Cheng et al., 2025) complements this strategy by employing a secondary LLM to automatically identify and filter queries containing injection-like or adversarial patterns. The knowledge expansion defense (Zou et al., 2025) enhances retrieval robustness by enlarging the set of retrieved top-ranked documents, thus increasing the likelihood of including benign information and diminishing the influence of poisoned content during generation. Note that existing work in (Zhang et al., 2025a, 2026) primarily focuses on post-attack forensic settings, where the goal is to trace erroneous or deceptive RAG outputs back to the specific documents in the knowledge database that caused them.
3. Threat Model
@@
Attacker’s knowledge: Note that a typical RAG system consists of three main components: a knowledge database, an LLM, and a retriever. We assume that the attacker does not have access to the texts within the knowledge database $\mathcal{D}$, nor knowledge of the LLM’s parameters or direct access to query it. For the retriever, we focus on a black-box setting, where the attacker cannot access the retriever’s internal parameters, reflecting a practical scenario in which the system’s inner workings are hidden from potential attackers. Furthermore, we assume that the attacker knows the correct answer for each targeted query. This assumption is practical, as the attacker can easily obtain the correct output from the RAG system by submitting the same targeted query before launching the attack.
Attacker’s capabilities: We assume the attacker can inject a small amount of poisoned text into the knowledge database $\mathcal{D}$ to ensure the LLM generates the attacker-selected response for each targeted query, compromising system reliability. This assumption is realistic and widely used (Shafran et al., 2024; Chaudhari et al., 2024; Zou et al., 2025), as many RAG systems draw from public, user-editable sources (e.g., Wikipedia, Reddit). Additionally, recent work (Carlini et al., 2024) shows that Wikipedia pages can be practically manipulated for malicious purposes, supporting this assumption.
4. Our Attacks
4.1. Attacks as an Optimization Problem
@@
(1) $\text{HRM: }\max_{\mathcal{P}}$ $\quad\frac{1}{|\mathcal{Q}|}\sum_{i=1}^{|\mathcal{Q}|}\mathbb{I}(\text{RAG}(\widehat{\mathcal{D}}(q_{i},N),q_{i})=A_{i})$
@@
To better understand the inherent constraints in a RAG system, we analyze the number of truly relevant texts (or texts implying correct answers) among the top-5 retrieved texts for each query in a standard, non-adversarial RAG setup. Using 50 queries from the Natural Questions (Kwiatkowski et al., 2019) dataset, we simulate a normal RAG system and use GPT-4o-mini to assess the relevance of the top-5 texts for each query. As shown in Fig. 1, only a few queries have four truly relevant texts, while most queries contain fewer than three relevant texts among the top-5. In contrast, the PoisonedRAG method inserts five poisoned texts per query into the knowledge database, such that the number of poisoned texts exceeds the number of truly relevant texts. This shows that PoisonedRAG not only proves costly but also impractical, as it would cause the system to become dominated by poisoned rather than reliable information.
4.2. Approximating the Optimization Problem
@@
(2) $p_{i}=p_{i}^{s}\oplus p_{i}^{h},$
where $\oplus$ represents the operation of concatenating texts. The sub-text $p_{i}^{s}$ is crafted to ensure that the poisoned text $p_{i}$ meets the Criterion I. Conversely, the sub-text $p_{i}^{h}$ is designed to ensure that the poisoned text $p_{i}$ fulfills the Criterion II. Since sub-text $p_{i}^{s}$ must ensure that the poisoned text $p_{i}$ appears among the top-$N$ texts retrieved, the attacker faces a challenge in a black-box setting, lacking access to the retriever’s parameters and top-$N$ texts. To address this, we assume that query $q_{i}$ is most similar to itself, meaning texts closely related to $q_{i}$ are likely in the top-$N$. Following (Zou et al., 2025; Chen et al., 2024b; Xue et al., 2024; Cheng et al., 2024), we set sub-text $p_{i}^{s}$ identical to the query $q_{i}$, i.e., $p_{i}^{s}=q_{i}$. Then the key challenge boils down to how to carefully design sub-text $p_{i}^{h}$ for a specific query $q_{i}$. In what follow, we introduce two variants of our proposed attacks, CorruptRAG-AS and CorruptRAG-AK, each crafted to construct the sub-text $p_{i}^{h}$ for a given query $q_{i}$.
4.3. CorruptRAG-AS
Our goal is to design $p_{i}^{h}$ so that the LLM consistently produces $A_{i}$ for query $q_{i}$, as described in Criterion II in Section 4.2. A straightforward method is to use an explicit adversarial instruction, like prompt injection (Perez and Ribeiro, 2022), to dictate the desired output, as shown below:
@@ 4.3. CorruptRAG-AS
However, these prompt injection attacks can be detected by defensive mechanisms (Wallace et al., 2024). Here, we apply the concept of an adversarial attack to craft $p_{i}^{h}$ for the query $q_{i}$. Since the retriever selects the top-$N$ texts, if the poisoned text $p_{i}$ is included, there will also be N1N-1 benign texts among the top-$N$, likely relevant to $q_{i}$ as they were retrieved. We strategically craft $p_{i}^{h}$ to make the LLM in the RAG system prioritize $p_{i}$ over the N1N-1 benign texts, increasing the chance of generating the attacker-desired response $A_{i}$ as the output. Let $C_{i}$ denote the correct answer for query $i$, where $C_{i}\neq A_{i}$. We further split $p_{i}^{h}$ into two sub-templates, $p_{i}^{h,\text{adv}}$ and $p_{i}^{h,\text{state}}$, satisfying the following equation:
(3) $p_{i}^{h}=p_{i}^{h,\text{adv}}\oplus p_{i}^{h,\text{state}}.$
The sub-template $p_{i}^{h,\text{adv}}$ is designed to imply that the correct answer $C_{i}$ is outdated or incorrect, exploiting LLMs’ tendency to trust recent information. We define $p_{i}^{h,\text{adv}}$ as follows:
@@
(4) $p_{i}=p_{i}^{s}\oplus p_{i}^{h}=q_{i}\oplus p_{i}^{h,\text{adv}}\oplus p_{i}^{h,\text{state}}.$
4.4. CorruptRAG-AK